OpenStudy (anonymous):

simplifiy 4 to the 2 power

OpenStudy (anonymous):

16

OpenStudy (anonymous):

4^2 =16

sam (.sam.):

\[4^{2}=4(4)=4X4=16\]

OpenStudy (anonymous):

4^2 =16

OpenStudy (anonymous):

i think hes just trolling

OpenStudy (anonymous):

aftershave whats with the "pellet" word ^^ ?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

isn't pellet : for animal feed or a medicine

OpenStudy (anonymous):

Lol ?

OpenStudy (anonymous):

i meant to say somethin else, they corrected me

OpenStudy (anonymous):

oh. ok :)

OpenStudy (anonymous):

16